### v1

@@ Abstract
Advances in AI portend a new era of sophisticated disinformation operations. While individual AI systems already create convincing—and at times misleading—information, an imminent development is the emergence of malicious AI swarms. These systems can coordinate covertly, infiltrate communities, evade traditional detectors, and run continuous A/B tests, with round-the-clock persistence. The result can include fabricated grassroots consensus, fragmented shared reality, mass harassment, voter micro-suppression or mobilization, contamination of AI training data, and erosion of institutional trust. With democratic processes worldwide increasingly vulnerable, we urge a three-pronged response: (1) platform-side defenses—always-on swarm-detection dashboards, pre-election high-fidelity swarm-simulation stress-tests, transparency audits, and optional client-side “AI shields” for users; (2) model-side safeguards—standardized persuasion-risk tests, provenance-authenticating passkeys, and watermarking; and (3) system-level oversight—a UN-backed AI Influence Observatory.
@@
Information operations long predate social media: Cold-War propaganda ranged from Soviet claims that AIDS was a U.S. bioweapon to documented covert U.S. efforts against Latin-American leaders. Online manipulation then accelerated in the 2000s–2010s, targeting Brexit, the 2016 U.S. election, and campaigns in Brazil and the Philippines, and even fueling violence such as the Rohingya genocide [11, 12, 13, 14, 15, 16]. Field experiments still find modest, if any, attitude shifts [17, 18], yet sustained state and private investment, along with documented links between hate campaigns and offline harm [19], ensure tactics will keep evolving.
@@
Second, agents now can map social graphs and slip into vulnerable communities with tailored appeals. Malicious agent swarm tactics include infiltrating vulnerable online communities to get humans to follow them and generating high volumes of content that is appealing to those communities [34]. AI agents can employ systems that map social graphs at scale, identify key communities and beliefs, and track trending topics [35]. This can also work in a decentralized fashion while still having global, network–wide efficacy [36]. Equipped with such capabilities, agent swarms can position themselves for maximum impact and tailor messages to the beliefs and cultural cues of each community, enabling more precise targeting than previous botnet approaches.
Third, human-level mimicry lets swarms evade detectors that once caught copy-paste bots. Methods to detect coordinated inauthentic behavior, such as influence operations, generally rely on activity patterns that are suspiciously similar across accounts and therefore statistically unlikely to result from independent behaviors [37, 38]. High-fidelity photorealistic avatars, context-aware slang, and heterogeneous posting rhythms circumvent the tell-tale synchrony that older detectors flag [39, 40].
Fourth, continuous A/B testing refines messages at machine speed. In the near term, human operators will still steer agent swarms while letting the agents perform limited in-context adaptation. Over time, however, the balance is likely to tilt toward self-optimizing swarms that harvest their own reward signals: real-time engagement data, structural cues from recommender systems, or even natural-language critiques [41]. Even these systems cannot improve without such measurable feedback. Given sufficient signals, though, swarms can run millions of micro-A/B tests, propagate the winning variants at machine speed, and iterate far faster than human campaigners.
@@
Figure 1: From Malicious AI Swarm to Democratic Harm. Concentric-circle schematic linking threat mechanics to civic outcomes. At the core, a Malicious AI Swarm supplies at least five key capabilities—decentralized orchestration, community infiltration, detection evasion, continuous optimization, and persistence (grey ring). These capabilities radiate outward into at least nine democratic-impact pathways, including synthetic consensus cascades, undermined collective intelligence, fragmented epistemic commons, LLM training-data poisoning, mass harassment, algorithmic overcompensation, FUD-induced disengagement, institutional legitimacy erosion, and engineered norm shifts and mass mobilization (outer ring). The upward arrow signals the causal flow from inner mechanics to systemic harm. Note that this flow is likely reciprocal, as impaired democratic functioning increases a society’s vulnerability to malicious AI swarms. FUD = Fear, Uncertainty, Doubt.
@@
Collaborating AI agents can tailor streams of misleading information to each sub-community’s linguistic, cultural, and emotional markers, weaving partially overlapping but strategically segmented realities. Unlike passive “filter bubbles,” these engineered realities are designed to keep groups apart except when not desired, making cross-cleavage consensus increasingly unfeasible. Once initiated, such streams can also spread via social contagion such that the effect of bots might cascade beyond the people to whom the bots are directly connected, to two or three degrees of separation, as is often empirically the case [47].
@@
The emergence of sophisticated, collaborative, and autonomous swarms of AI agents capable of orchestrating influence operations marks a critical juncture. The causal arrow runs both ways: while malicious swarms endanger democratic norms, the quality of democratic governance will in turn shape how potent, or containable, those swarms become. A full shift to continuous detection, simulation, and proactive, defense-oriented AI must therefore be implemented and operational before AI swarm capabilities reach mass availability.
@@
Finally, frontier models must undergo standardized persuasion-risk evaluations and publish scores in model cards before release. Companies should disclose within 24 hours whether entities interacting with users are human or automated. Pre-bunking campaigns that educate citizens about AI tactics can build cognitive immunity, but paradoxically, the very alertness they foster can still be exploited by threat actors to deepen FUD. “Prosocial media” architectures on interoperable platforms can offer structural resilience [57]. Governments and tech firms must prioritize AI-safety R&D and fund independent measurement of misuse potential and societal impact.
The next years give democracies a fleeting, actionable window to pull AI-enabled influence operations back from the brink. If platforms deploy always-on swarm detectors, frontier labs submit models to standardized persuasion “stress-tests,” and governments launch an AI Influence Observatory that publishes open incident telemetry, we can blunt the most destabilizing tactics before critical political future events, without freezing innovation. Doing so demands the same choreography already mastered by AI swarms: rapid iteration, transparent data-sharing, and tight coordination across scientific, civil-society, industry, and election-security teams. Like earlier calls to confront emerging bio-risks, success hinges on acting “collaboratively … without unnecessarily impeding scientific research,” while keeping the public square resilient and accountable [58]. By committing now to rigorous measurement, proportionate safeguards, and shared oversight, upcoming elections could even become a proving ground—not a casualty—of principled, democratic AI governance.

### v2

@@ Abstract
Advances in AI portend a new era of sophisticated disinformation operations. While individual AI systems already create convincing—and at times misleading—information, an imminent development is the emergence of malicious AI swarms. These systems can coordinate covertly, infiltrate communities, evade traditional detectors, and run continuous A/B tests, with round-the-clock persistence. The result can include fabricated grassroots consensus, fragmented shared reality, mass harassment, voter micro-suppression or mobilization, contamination of AI training data, and erosion of institutional trust. With increasing vulnerabilities in democratic processes worldwide, we urge a three-pronged response: (1) platform-side defenses—always-on swarm-detection dashboards, pre-election high-fidelity swarm-simulation stress-tests, transparency audits, and optional client-side “AI shields” for users; (2) model-side safeguards—standardized persuasion-risk tests, provenance-authenticating passkeys, and watermarking; and (3) system-level oversight—a UN-backed AI Influence Observatory.
@@
Information operations long predate social media: Cold-War propaganda ranged from Soviet claims that AIDS was a US bioweapon to documented covert U.S. efforts against Latin-American leaders. Online manipulation then accelerated in the 2000s–2010s, targeting Brexit, the 2016 US election, and campaigns in Brazil and the Philippines, and even fueling violence such as the Rohingya genocide [11, 12, 13, 14, 15, 16]. Field experiments still find modest, if any, attitude shifts [17, 18], yet sustained state and private investment, along with documented links between hate campaigns and offline harm [19], ensure tactics will keep evolving.
@@
Second, agents now can map social graphs and slip into vulnerable communities with tailored appeals. Malicious agent swarm tactics include infiltrating vulnerable online communities to get humans to follow them and generating high volumes of content that appeals to those communities [34]. AI agents can employ systems that map social graphs at scale, identify key communities and beliefs, and track trending topics [35]. This can also work in a decentralized fashion while still having global, network–wide efficacy [36]. Equipped with such capabilities, agent swarms can position themselves for maximum impact and tailor messages to the beliefs and cultural cues of each community, enabling more precise targeting than previous botnet approaches.
Third, human-level mimicry lets swarms evade detectors that once caught copy-paste bots. Methods for detecting coordinated inauthentic behavior, such as influence operations, generally rely on activity patterns that are suspiciously similar between accounts and therefore statistically unlikely to result from independent behaviors [37, 38]. High-fidelity photorealistic avatars, context-aware slang, and heterogeneous posting rhythms circumvent the tell-tale synchrony that older detectors flag [39, 40].
Fourth, continuous A/B testing refines messages at machine speed. In the near term, human operators will still steer agent swarms while letting the agents perform limited in-context adaptation. However, over time, the balance is likely to tilt toward self-optimizing swarms that harvest their own reward signals: real-time engagement data, structural cues from recommender systems, or even natural-language critiques [41]. Even these systems cannot improve without such measurable feedback. Given sufficient signals, though, swarms can run millions of micro-A/B tests, propagate the winning variants at machine speed, and iterate far faster than human campaigners.
@@
Figure 1: From Malicious AI Swarm to Democratic Harm. Concentric-circle schematic linking threat mechanics to civic outcomes. At the core, a Malicious AI Swarm supplies at least five key capabilities—decentralized orchestration, community infiltration, detection evasion, continuous optimization, and persistence (inner ring). These capabilities radiate outward into at least nine democratic-impact pathways, including synthetic consensus cascades, undermined collective intelligence, fragmented epistemic commons, LLM training-data poisoning, mass harassment, algorithmic overcompensation, FUD-induced disengagement, institutional legitimacy erosion, and engineered norm shifts and mass mobilization (outer ring). The upward arrow signals the causal flow from inner mechanics to systemic harm. Note that this flow is likely reciprocal, as impaired democratic functioning increases a society’s vulnerability to malicious AI swarms. FUD = Fear, Uncertainty, Doubt.
@@
Collaborating AI agents can tailor streams of misleading information to each sub-community’s linguistic, cultural, and emotional markers, weaving partially overlapping but strategically segmented realities. Unlike passive “filter bubbles,” these engineered realities are designed to keep groups apart except when not desired, making cross-cleavage consensus increasingly unfeasible. Once initiated, such streams can also spread via social contagion, with the effect of bots potentially cascading beyond the people to whom the bots are directly connected, to two or three degrees of separation, as is often empirically the case [47].
@@
The emergence of sophisticated, collaborative, and increasingly autonomous swarms of AI agents capable of orchestrating influence operations marks a critical juncture. The causal arrow runs both ways: while malicious swarms endanger democratic norms, the quality of democratic governance will in turn shape how potent, or containable, those swarms become. A full shift to continuous detection, simulation, and proactive, defense-oriented AI must therefore be implemented and operational before AI swarm capabilities reach mass availability.
@@
Finally, frontier models must undergo standardized persuasion-risk evaluations and publish scores in model cards before release. Companies should disclose within 24 hours whether entities interacting with users are human or automated. Pre-bunking campaigns that educate citizens about AI tactics can build cognitive immunity, but paradoxically, the very alertness they foster can still be exploited by threat actors to deepen FUD. “Prosocial media” architectures on interoperable platforms can offer structural resilience [57]. Governments and tech firms must prioritize AI-safety R&D while funding independent assessments of misuse potential and societal impact.
In the coming years, democracies have a brief but crucial opportunity to pull AI-enabled influence operations back from the brink. If platforms deploy always-on swarm detectors, frontier labs submit models to standardized persuasion stress tests, and governments launch an AI Influence Observatory that publishes open incident telemetry, we can blunt the most destabilizing tactics before critical political future events, without freezing innovation. Doing so demands the same choreography already mastered by AI swarms: rapid iteration, transparent data-sharing, and tight coordination across scientific, civil-society, industry, and election-security teams. Like earlier calls to confront emerging bio-risks, success hinges on acting “collaboratively … without unnecessarily impeding scientific research,” while keeping the public square resilient and accountable [58]. By committing now to principled action, upcoming societal decisions could even become a proving ground—rather than a casualty—of democratic AI governance.
Acknowledgements
We would like to thank Jon Roozenbeek and Asbjørn Følstad for their valuable input to this paper. We are also grateful to Patrycja Antosz, Önder Gürcan, Ivan Puga Gonzalez, and Phil Tinn for the fruitful discussions that contributed to the development of this work.